Hello:
I am trying to find out how to work this question:
If Sally can paint a house in 4 hours and John can paint the same house in six hours, how long will it take for the both of them to paint the house together?
The answer says it's 2 Hours 24 Minutes
I can not, for the life of me, figure out how they got this answer. I have tried to work it multiple ways, but can't come up with that answer. PLEASE HELP!!!

Answer by stanbon(75887)

If Sally can paint a house in 4 hours and John can paint the same house in six hours, how long will it take for the both of them to paint the house together?
-----------------------
Sally DATA;
time = 4 hr/job ; rate = 1/4 job/hr
--------------------------------------
John DATA:
time = 6 hr/job ; rate = 1/6 job/hr
----------------------------------------
Together DATA:
time = x hr/job ; rate = 1/x job/hr
-------------------------------------------
Equation:
rate + rate + together rate
1/4 + 1/6 = 1/x
Multiply thru by 12x to get:
3x + 2x = 12
5x = 12
x = 2.4 hrs (time to do the job together)
Note: 0.4 hours = 0.4*60 minutes = 24 minutes
-----------------------------
2.4 hrs = 2 hrs 24 minutes
================================
Cheers,
Stan H.